# Self-Enhancing Video Data Management System for Compositional Events with Large Language Models

### v1

@@ 4. Proposing and generating new UDFs
VOCAL-UDF needs to address the following challenges. First, it must determine whether existing UDFs can adequately parse a user query, or if new UDFs must be created (C1). Second, VOCAL-UDF should have the capability to implement UDFs across a diverse range of semantic concepts (C2). Third, since VOCAL-UDF utilizes LLMs to generate UDFs, which can be prone to errors, it is crucial to maintain high quality in the final UDFs it produces (C3). We now discuss our solutions in the following subsections. Section 4.1 presents how VOCAL-UDF translates queries and proposes new UDFs. Section 4.2 and Section 4.3 discuss how VOCAL-UDF supports both program-based UDFs and distilled-model UDFs to manage diverse semantic concepts. Finally, we discuss UDF selection in Section 4.4.
@@ 4.1. Query parsing and UDF proposal
The Query Parser is responsible for converting natural language queries into our DSL format and determining if new UDFs need to be created. User queries can range from detailed and explicit descriptions to highly subjective ones (e.g., a dangerous traffic situation). Conversely, the resulting DSL queries must adhere strictly to the DSL grammar and use only the existing UDFs. VOCAL-UDF must understand the semantics of both the user’s query and the available UDFs, mapping each part of the query to an existing UDF or suggesting the creation of a new one. Moreover, VOCAL-UDF needs to be resilient to the linguistic ambiguities and synonymous terms in natural language queries.
To convert from a rich and diverse natural language to a consistent DSL format, VOCAL-UDF utilizes LLMs. LLMs have shown strong capabilities in SQL and program generations (Pourreza and Rafiei, 2023; Wang et al., 2023a; Gupta and Kembhavi, 2023) using only in-context learning without fine-tuning. They excel at understanding and performing natural language tasks due to their training on vast datasets. To ensure generated DSL queries strictly follow the DSL grammar, we incorporate domain-specific constraints when prompting LLMs and conduct post-verification to ensure syntax correctness.
Inspired by Wang et. al. (Wang et al., 2023a) whose method provides a formal grammar during in-context learning for DSL generation and Hsieh et. al. (Hsieh et al., 2023a) whose approach showcases LLMs’ potential for zero-shot tool-usage through tool documentation, VOCAL-UDF combines both prompting techniques to provide the DSL definition, UDF format, and descriptions of available UDFs as natural language documentations. This enables VOCAL-UDF to generate DSL queries that adhere to the grammar, determine whether new UDFs need to be created, and interact with LLMs in a zero-shot manner. Figuring out the optimal prompting strategy is not the focus of the paper. An example prompt is shown in LABEL:fig:parsing_query_prompt and LABEL:fig:proposing_udf_prompt.
@@
To enhance the reliability of LLM’s response, VOCAL-UDF performs a post-verification of the generated DSL query. Specifically, VOCAL-UDF interprets the DSL string to ensure it follows the grammar rules and all predicates are resolved by the available UDFs. If parse errors are encountered, VOCAL-UDF appends the error message to the context and asks the LLM to make another attempt.
@@
When the query contains predicates that cannot be resolved using the available UDFs, the LLM identifies this and proposes new UDFs. The output from this process is the function signature and textual description of the proposed UDF generated by the LLM. In case of a query with multiple missing UDFs, VOCAL-UDF returns a list of proposed UDFs and will generate them one by one. LABEL:lst:proposed_udf shows an example proposed UDF of behind.
@@ 4.2. Program-based UDF generation
The task of the UDF Generator is to implement executable UDFs based on the proposed function’s type signatures and descriptions (LABEL:lst:proposed_udf) generated by the LLM. As discussed, VOCAL-UDF should produce high-quality UDFs (C3) for a wide range of semantic concepts (C2). VOCAL-UDF, again, leverages the programming capabilities of LLMs (Rozière et al., 2023; Bubeck et al., 2023; Chen et al., 2023b, 2022b) to implement proposed UDFs as Python programs. However, this does not solve all the problems. First, some concepts require understanding the visual content of the image, such as determining whether the make of a car is Subaru, which is difficult to solve by writing Python programs. Second, even the most advanced LLMs today are still error-prone, and their programming performance in complex and nuanced tasks is still not on par with humans.
To address the first problem, VOCAL-UDF supports two types of UDF implementations: Program-based UDFs, which are Python programs generated by LLMs (Section 4.2), and distilled-model UDFs, which are lightweight vision models distilled from strong pretrained models (Section 4.3). To improve generated UDF quality, VOCAL-UDF uses a two-step approach by verifying both syntax correctness (Section 4.2.2) and semantic correctness (Section 4.4).
Given a UDF signature and description , a video database instance over schema , VOCAL-UDF performs the following steps to generate a UDF :
@@
UDF generation with LLMs: Given , , , and , VOCAL-UDF prompts an LLM to generate a list of candidate Python functions .
@@
Syntax verification (only for program-based UDFs): For each candidate program , VOCAL-UDF executes it on a small set of data sampled from to verify the syntax correctness, correct errors, and filter out invalid programs after a few trials.
@@
Semantic verification: VOCAL-UDF evaluates the semantic correctness of the remaining candidates and selects the best one (Section 4.4).
We describe each step in detail below.
4.2.1. Program generation with LLM
In its most basic form, VOCAL-UDF prompts the LLM to generate a program based on the UDF signature and description . To ensure can be integrated as a UDF in SQL, VOCAL-UDF rewrites in the DSL format to that accepts columns from the schema as inputs. For example, the behind function is rewritten from behind(o0, o1) to behind(o1_y1, o1_y2, o2_y1, o2_y2), where o1_y1, o1_y2, o2_y1, o2_y2 are columns from . The generated Python program’s input is a set of attribute values that correspond to one or two objects in the video database, depending on the number of variables in . It generates code that operates on those values, including the pixels column, and returns a boolean indicating if the predicate is satisfied or not. Figure 3 (program ) shows an example program UDF for the behind() predicate.
To generate UDFs with good quality, VOCAL-UDF must resolve three challenges. First, natural language descriptions are often ambiguous and may not capture all the details of a user’s intent (Lahiri et al., 2023). For example, one user can interpret the “far” relationship between two objects as a distance greater than 10 meters, while another user may interpret it as a distance greater than three times the size of the larger object. Second, predicates in video compositional queries often include parameters that need to be tuned (Mell et al., 2023, 2024) for different datasets and user intents, e.g., determining the threshold distance for the “far” relationship, or the color range for the “color-silver” attribute. Lastly, as VOCAL-UDF expands its collection of UDFs and database incrementally, it is essential that new UDFs be able to utilize results from previously established ones.
To address the challenge of linguistic ambiguity, instead of asking the LLM to generate one program , we provide the LLM with and and ask it to generate a list of candidate programs with different semantic interpretations to cover a wide range of possible implementations: , where and denote the Python program and the semantic interpretation of the -th candidate. LABEL:fig:implementing_udf_prompt shows an example prompt used by VOCAL-UDF in this step. Later, VOCAL-UDF will verify and select one program for each proposed UDF. Figure 3 shows three example candidate programs, each with a different semantic interpretation.
@@ 4.2.1. Program generation with LLM
Overall, program-based UDFs are well-suited for concepts involving bounding box-like spatial relationships. In addition, other attributes as well as the pixels column can also be used to reason about existing concepts and perform statistical analysis of frames pixels. With frame pixels, a program-based UDF can also invoke pretrained models, but they are expensive and slow in general. As discussed in Section 4.3, distilled-model UDF is a more efficient approach for concepts that require visual understanding of videos.
@@ 4.2.2. Syntax verification
Prior works have proposed various approaches to improve the performance of LLM-generated programs: The first line of work uses unit tests to verify the functional correctness of generated programs (Cai et al., 2023; Hu et al., 2023; Chen et al., 2023c, 2021). This approach is not suitable for VOCAL-UDF because users do not know in advance what UDFs will be generates, and thus cannot provide labeled data before issuing a query. Another line of work leverages LLMs to automatically generate test cases (Chen et al., 2022b), evaluate the generated programs (Chen et al., 2023b; Shinn et al., 2023), and select the best one (Wang et al., 2023b). However, the ambiguity of semantic concepts in video queries means that the correct program is not always unique and often many of the generated programs are reasonable. As a result, the best program cannot be easily identified without user feedback and a working dataset.
VOCAL-UDF uses a two-step approach to verify and select the best program: syntax verification and semantic verification. In the first step, VOCAL-UDF focuses solely on the syntax correctness of the programs. VOCAL-UDF executes each candidate program on a small sample of data from the database and checks: (i) whether the number and types of inputs and outputs are correct, (ii) whether the program can be executed with the data samples, and (iii) whether , if any, can be parsed successfully. If the verification fails, VOCAL-UDF appends the error message to the context and prompt the LLM to make another attempt. A sample is a tuple that contains attributes values corresponding to two objects from the database if the target concept is a relationship, or to one object if the target concept is an attribute. If the program still fails after a few trials, VOCAL-UDF discards the program. In our prototype, we empirically set the number of trials to five. In the second step, VOCAL-UDF utilizes user labels to select the program that best align with user intent, which we discuss in Section 4.4.
@@ 4.3. Distilled-model UDF generation
As discussed, while program-based UDFs are powerful and flexible for predicates that reason about existing concepts, bounding box coordinates, or perform a simple statistical analysis of pixel values, they struggle with tasks that require understanding the visual contents of frames. Even though pretrained models like VLMs can be used in a program-based UDF to classify relationships and attributes in a zero-shot manner, running such models over the entire video dataset is prohibitively expensive. For instance, the estimated expense of sequentially processing the CLEVRER (Yi et al., 2020) dataset of 10,000 videos with GPT-4V for the “color-red” attribute amounts to $2,826 and would take 24 years. The goal of VOCAL-UDF is to generate cheaper UDFs with visual understanding capabilities. However, training a lightweight image classifier from scratch would require the user to spend a lot of time labeling for just one concept.
Model distillation is a common technique in machine learning to transfer knowledge from a large model to a smaller, more efficient model (Hsieh et al., 2023b; Toubal et al., 2024; Bucila et al., 2006; Chen et al., 2017; Hinton et al., 2015). Modeling Collaborator (Toubal et al., 2024) is a newly proposed framework that leverages foundation models to train image classifiers for visual concepts using minimal user effort and it goes through the following steps. Given a target concept and description, the system (i) mines relevant images from the public domain, (ii) uses a mixture of foundation models to annotate sampled images, (iii) trains a lightweight image classifier using image features extracted from a pretrained model (e.g., CLIP) and image labels annotated by the foundation models, and (iv) performs multiple rounds of active learning to further improve its performance.
In VOCAL-UDF, we adopt a similar approach to automatically construct lightweight image classifiers for new relationships and attributes without any user labeling effort, but with the necessary modifications to resolve three unique challenges presented in our compositional query setting. First, random sampling of the user dataset might not give enough positive samples for training, especially when the target concept is less common. Second, VOCAL-UDF generates UDFs for relationships and attributes, which differ from the concepts in (Toubal et al., 2024). An image usually includes multiple objects, and VOCAL-UDF needs a different prompting strategy to guide VLMs in classifying specific objects or pairs of objects. Finally, for the same reason, simply extracting features from the entire image is insufficient for training a good classifier.
@@ 4.3.1. Image sampling
By default, to generate a distilled-model UDF, VOCAL-UDF randomly samples frames from the user’s video dataset for annotation. However, when the target visual concept is infrequent in the dataset, random sampling does not effectively collect enough positive samples for training. For instance, only 0.83% human-object pairs have an “eating” relationship in the Charades (Ji et al., 2020) dataset. Our solution is object-aware sampling to bootstrap the sampling process. Since all objects of interest are already detected and tracked in the video dataset, VOCAL-UDF can filter out objects that are not relevant to the target concept and only sample objects that can possibly be involved in the concept. For example, when labeling the “eating” relationship, object classes like “food” and “person” are more relevant than “car” and “window”. To do this, VOCAL-UDF first asks an LLM for relevant object classes, and then only samples objects belonging to these classes.

### v2

@@ 4. VOCAL-UDF approach
VOCAL-UDF needs to address several challenges. First, it must determine whether existing UDFs can adequately answer a user query, or if new UDFs should be created (C1). Second, VOCAL-UDF should support implementing UDFs drawn from diverse range of semantic concepts (C2). Third, since VOCAL-UDF utilizes error-prone LLMs to generate UDFs, it is crucial to ensure high quality in the produced UDFs (C3). We now discuss our solutions to each challenge.
Figure 2 shows the architecture of VOCAL-UDF. The user initializes the system with a video dataset and an optional set of UDFs. After preprocessing (Section 5), the user can issue NL queries to identify compositional events within the videos. The Query Parser (Section 4.1) parses the query into the DSL notation described in Section 2. If successful, the DSL query is passed to the Query Executor to find all matching videos in the dataset. If the query contains predicates that existing UDFs cannot resolve, the UDF Proposer (Section 4.1) is invoked to propose the names and descriptions of new UDFs. The UDF Generator (Sections 4.2 and 4.3) creates executable candidates for each proposed UDF. Next, the UDF Selector (Section 4.4) solicits user labels to select the implementation that best aligns with the user’s intent. With updated UDFs, the Query Parser re-parses the query for execution. The Storage Manager maintains the raw videos, extracted relational information, and available UDFs.
@@ 4.1. Query parsing and UDF proposal
The Query Parser converts NL queries into our DSL and determines the need for new UDFs. VOCAL-UDF must understand the semantics of the user’s query and available UDFs, mapping each part of the query to an existing UDF or suggesting the creation of a new one. Moreover, VOCAL-UDF needs to be resilient to the linguistic ambiguities and synonymous terms in NL queries.
VOCAL-UDF utilizes LLMs to convert complex NL to a consistent DSL format. LLMs show strong capabilities in SQL and program generations (Pourreza and Rafiei, 2023; Wang et al., 2023a; Gupta and Kembhavi, 2023) via in-context learning without fine-tuning. We incorporate domain-specific constraints into LLM prompts to ensure adherence to DSL grammar and conduct post-verification to ensure syntax correctness. Inspired by Wang et. al. (Wang et al., 2023a) and Hsieh et. al. (Hsieh et al., 2023a), VOCAL-UDF provides the DSL definition, UDF format, and descriptions of available UDFs as NL documentation. This enables VOCAL-UDF to generate grammatically correct DSL queries, determine whether new UDFs need to be created, and interact with LLMs in a zero-shot manner (C1). While VOCAL-UDF focuses on zero-shot prompting, introducing few-shot examples (Brown et al., 2020) could potentially further improve UDF proposal performance. However, referencing unavailable UDFs in examples may lead LLMs to incorrectly assume their availability at query time. Since the available UDFs can change as the database evolves, examples must be carefully curated for different database states. Figuring out the optimal prompting strategy, though beneficial, is not the focus of this paper. An example prompt is shown in LABEL:fig:parsing_query_prompt and LABEL:fig:proposing_udf_prompt.
@@
To enhance the reliability of the LLM’s response, VOCAL-UDF performs post-verification of the generated DSL query to ensure it is syntactically valid and only uses available UDFs. If parse errors are encountered, VOCAL-UDF appends the error message to the context and asks the LLM to make another attempt.
@@
When the query contains predicates that cannot be resolved using the available UDFs, we prompt the LLM to identify this and propose new UDFs. The output from this process is the function signature and textual description of the proposed UDF generated by the LLM. In case of a query with multiple missing UDFs, VOCAL-UDF returns a list of proposed UDFs and will generate them one by one. An example proposed UDF of behind is:
Since both the Query Parser and UDF Proposer utilize LLMs, their outputs may sometimes be inaccurate. If the LLM proposes fewer UDFs than needed, it may make performance improvements less noticeable but will not degrade performance. Conversely, proposing more UDFs can increase system runtime and monetary costs; however, as detailed in Section 4.4, we have implemented techniques to ensure that generated UDFs (whether superfluous or inaccurate) do not adversely affect performance. Finally, users may also rephrase their NL query if they are not fully satisfied with the query results.
We show the effectiveness of our approach empirically in Section 6 for queries with detailed and explicit descriptions. Semantic parsing (e.g., text-to-SQL) is a highly active research area with significant recent advances (Katsogiannis-Meimarakis and Koutrika, 2023; Deng et al., 2022; Zhong et al., 2017; Yu et al., 2018; Gao et al., 2024). While current methods still fall short of human-level performance (Li et al., 2023a), improving text-to-SQL is not the goal of our paper. VOCAL-UDF focuses on the efficient generation of missing UDFs, relies on LLMs currently adequate performance for text-to-SQL in our context, and will benefit from future advances in this area.
@@ 4.2. Program-based UDF generation
The UDF Generator implements executable UDFs based on the LLM-proposed UDF signatures and descriptions. As discussed, VOCAL-UDF should produce high-quality UDFs (C3) for a wide range of semantic concepts (C2). While VOCAL-UDF leverages the programming capabilities of LLMs (Rozière et al., 2024; Bubeck et al., 2023; Chen et al., 2024b, 2023) to generate UDFs as Python programs, more work is required. Some tasks require complex visual understanding, rendering it challenging to solve via programs (e.g., determining a car’s make). Additionally, even the most advanced LLMs remain error-prone and their programming performance for nuanced tasks is not yet on par with humans.
To address the first problem, VOCAL-UDF supports two types of UDF implementations: program-based UDFs, which are Python programs generated by LLMs (Section 4.2), and distilled-model UDFs, which are lightweight vision models distilled from strong pretrained models (Section 4.3). To improve generated UDF quality, VOCAL-UDF uses a two-step approach by verifying both syntactic (Section 4.2.2) and semantic (Section 4.4) correctness.
At a high level, given a UDF signature and description , a video database instance over schema , VOCAL-UDF performs the following steps to generate a UDF :
@@
Generate UDF candidates using an LLM. Given , , , and , VOCAL-UDF prompts an LLM to generate a set of candidate Python functions .
@@
Syntactically verify candidates. VOCAL-UDF executes each candidate program on a small dataset sampled from to verify syntactic correctness, ensuring that all passing candidates are executable.
@@
Semantically verify candidates. VOCAL-UDF finally evaluates the semantic correctness of the remaining candidates and selects the best one (see Section 4.4).
We next describe each step in further detail.
4.2.1. Generating candidate programs using an LLM
In its most basic form, VOCAL-UDF prompts the LLM to generate a program based on the UDF signature and description . To ensure can be eventually expressed as a SQL predicate, VOCAL-UDF rewrites in the DSL format to that accepts columns from the schema as inputs. For example, the behind function is rewritten from behind(o0, o1) to behind(o1_y1, o1_y2, o2_y1, o2_y2), where o1_y1, o1_y2, o2_y1, o2_y2 are columns of . The generated Python program’s input is a set of attribute values that correspond to one or two objects in the video database, depending on the number of variables in . It generates code that operates on those values, including the pixels column, and returns a boolean indicating if the predicate is satisfied or not. Figure 3 (program ) shows an example program UDF for the behind() predicate.
To generate quality UDFs, VOCAL-UDF must resolve three challenges. First, NL descriptions are often ambiguous and may not entirely capture the user’s intent (Lahiri et al., 2023) (e.g., user-specific definition of “far”). Second, predicates in video compositional queries often include hyperparameters that need to be tuned (Mell et al., 2023, 2024) for different datasets and user intents, e.g., determining the threshold distance for the “far” relationship. Lastly, as VOCAL-UDF expands its collection of UDFs and database incrementally, it is essential that new UDFs be able to utilize results from previously established ones.
To address the challenge of linguistic ambiguity, we provide the LLM with and and ask it to generate a list of candidate programs with a variety of semantic interpretations: , where and respectively denote a Python program and its semantic interpretation. LABEL:fig:implementing_udf_prompt shows an example prompt used by VOCAL-UDF in this step. Later, VOCAL-UDF verifies and selects one program for each proposed UDF. See Figure 3 for candidate program examples.
@@ 4.2.1. Generating candidate programs using an LLM
Overall, program-based UDFs are well-suited for concepts involving bounding box-like spatial relationships. In addition, other attributes and the pixels column can also be used to reason about existing concepts and perform statistical analysis of the pixels in a frame. A program-based UDF could also invoke a pretrained model over each frame, but this is in general expensive and slow. As discussed in Section 4.3, a distilled-model UDF is a more efficient approach for concepts that require visual understanding of videos.
@@ 4.2.2. Syntax verification
Prior works have proposed various approaches to improve the performance of LLM-generated programs. One line of work leverages unit tests to verify the functional correctness of generated programs (Cai et al., 2024; Hu et al., 2023; Chen et al., 2024c, 2021). This approach is not suitable for VOCAL-UDF because users do not know in advance what UDFs will be generates, and thus cannot provide labeled data before issuing a query. Another line of work leverages LLMs to automatically generate test cases (Chen et al., 2023), evaluate the generated programs (Chen et al., 2024b; Shinn et al., 2023), and select the best one (Wang et al., 2023b). However, the ambiguity of semantic concepts in video queries means that the correct program is not always unique and often many of the generated programs are reasonable. As a result, the best program cannot be easily identified without user feedback and a working dataset.
Additionally, the quality of an LLM is not guaranteed. To address this issue, VOCAL-UDF uses a two-step approach to verify and select the best program: syntax verification and semantic verification. In the first step, VOCAL-UDF focuses solely on the syntax correctness of the programs. VOCAL-UDF executes each candidate program on a small sample of data from the database and checks whether: (i) the number and types of inputs and outputs are correct, (ii) the program can be executed with the data samples, and (iii) , if any, can be parsed successfully. If the verification fails, VOCAL-UDF appends the error message to the context and prompt the LLM to make another attempt. VOCAL-UDF draws samples by constructing tuples that contain attributes and values that respectively correspond to one or two objects in the database. If the program still fails after a few trials, VOCAL-UDF discards the program. In our prototype, we empirically set the number of trials to five. Recent research has demonstrated the potential for formally verifying the correctness of generated UDFs based on a given semantic interpretation (Chen et al., 2025), which could be incorporated into VOCAL-UDF to further enhance the reliability of its generated Python programs. After generating multiple candidate programs, VOCAL-UDF next utilizes user labels to select the one that best aligns with user intent, which we discuss in Section 4.4.
@@ 4.3. Distilled-model UDF generation
While program-based UDFs are powerful and flexible for predicates that reason about existing concepts, bounding box coordinates, or perform a simple statistical analysis of pixel values, they struggle with tasks that require understanding the visual contents of frames. Even though pretrained models like VLMs can be used in a program-based UDF to classify relationships and attributes in a zero-shot manner, running such models over the entire video dataset is expensive. For instance, applying GPT-4o to every frame of the 10,000 five-second videos in the CLEVRER dataset (Yi et al., 2020) to evaluate a predicate (e.g., “color-red” or some more complex predicate) would cost approximately $1,413. This cost remains high even when applying techniques such as downsampling or predicate push-down: There are about 1.28M video frames and 5.44M objects to consider. Even if 99% of the image patches corresponding to these objects were filtered, the cost of evaluating a predicate on the remaining 1% of objects would still be around $14—far higher than generating one UDF (about $0.15; see Section 6.1). VOCAL-UDF’s goal is to generate cheaper UDFs with visual understanding capabilities. However, training a lightweight image classifier from scratch would require the user to spend a lot of time labeling for just one concept.
Model distillation is a common technique in machine learning to transfer knowledge from a large model to a smaller, more efficient model (Hsieh et al., 2023b; Toubal et al., 2024; Bucila et al., 2006; Chen et al., 2017; Hinton et al., 2015). Modeling Collaborator (Toubal et al., 2024) is a newly proposed framework that leverages foundation models to train image classifiers for visual concepts using minimal user effort. To do so, given a target concept and description, the system (i) mines relevant images from the public domain, (ii) uses foundation models to annotate sampled images, (iii) trains a lightweight classifier using features extracted from a pretrained model (e.g., CLIP) and labels annotated by the foundation models, and (iv) performs multiple rounds of active learning to further improve its performance.
In VOCAL-UDF, we adopt a similar approach to automatically construct lightweight image classifiers for new concepts without requiring user labeling but with the modifications needed to resolve three unique challenges presented in our compositional query setting. First, random sampling of the user dataset might not give enough positive samples for training, especially for rare concepts. Second, VOCAL-UDF generates UDFs for relationships and attributes, which differ from the concepts in (Toubal et al., 2024). An image usually includes multiple objects, and VOCAL-UDF needs a different prompting strategy to guide a VLM in classifying specific objects or pairs of objects. Finally, for the same reason, merely extracting features from the entire image is insufficient to train a good classifier.
@@ 4.3.1. Image sampling
To generate a distilled-model UDF, VOCAL-UDF initially randomly samples frames from the user’s video dataset for annotation. However, when the target visual concept is infrequent in the dataset, random sampling does not effectively collect enough positive samples for training. For instance, only 0.83% human-object pairs have an “eating” relationship in the Charades (Ji et al., 2020) dataset. Our solution is to apply object-aware sampling to bootstrap the sampling process. Since all objects of interest are already detected and tracked in the dataset, VOCAL-UDF can filter out irrelevant objects and sample only those likely to be involved in the target concept. For example, when labeling the “eating” relationship, object classes like “food” and “person” are more relevant than “car” and “window”. To do this, VOCAL-UDF first asks an LLM for relevant object classes, and then only samples objects belonging to these classes. LABEL:fig:filtering_objects_prompt shows an example prompt of object-aware sampling.